OpenStudy (anonymous):

Find the x- and y- intercepts -3x+8y=24

OpenStudy (anonymous):

Okay for y put x=0 and vice-versa

OpenStudy (anonymous):

x intercept -8 y intercept 3

OpenStudy (anonymous):

-3(0)-8y=24 -8y=24 - 8y/8=24/8 y=-3

OpenStudy (anonymous):

8(0)-3x=24 -3x=24 -3x/3=-24/3 x=-8

OpenStudy (anonymous):

x-intercept is correct and y-intercept is wrong its not -3(0)-8y=24 its -3(0)+8y=24

OpenStudy (anonymous):

but the equation is -3y +8x =24@loraaine

OpenStudy (anonymous):

-3(0)+8y=24 - 8y=24 -8y/8=24/8 y=-3

OpenStudy (anonymous):

how you got negative 8y?????????

OpenStudy (anonymous):

ok who is messing my fragile mind up ??

OpenStudy (anonymous):

-3x+8y=24 original question

OpenStudy (anonymous):

so -3x0 + 8y =24 +8y=24 y =3

OpenStudy (anonymous):

ok so now my x-y-intercepts are (0,3),(-8,0) correct?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

: ) Thank you